OpenStudy (anonymous):

set up all six possible triple integrals of the tetrahedron enclosed by the coordinate planes and the plane x + 2y + 3z = 30

OpenStudy (tkhunny):

Divide by 30, creating the "Intercept Form", and it should be relatively easy. \(\dfrac{x}{30} + \dfrac{y}{15} + \dfrac{z}{10} = 1\)

OpenStudy (anonymous):

im trying to post a picture of my paper but apparently its not working

OpenStudy (anonymous):

\[\int\limits_{0}^{10} \int\limits_{0}^{3z + 15} \int\limits_{0}^{2y -30} dx dy dz\]

OpenStudy (anonymous):

\[\int\limits_{0}^{15} \int\limits_{0}^{(2/3)y + 10} \int\limits_{0}^{(z-10)/3} dx dz dy\]

OpenStudy (anonymous):

check the first two please!

OpenStudy (anonymous):

\[\int\limits_{0}^{30} \int\limits_{0}^{1/2(x) +15} \int\limits_{0}^{2/3 (y) + 10} dz dy dx\]

OpenStudy (tkhunny):

No good. The most interior limits must have two variables, x:[0,30 - 2y - 3z] y: [0,15 - (3/2)z] z: [0,10] x:[0,30 - 2y - 3z] z: [0,10 - (2/3)z] y: [0,15] etc.